OpenStudy (aaronandyson):

2 cosec^2 30 + x sin^60 - tan^2 30 = 10 find x

OpenStudy (aaronandyson):

@triciaal @DivineSolar @danish071996 @dan815 @ganeshie8

OpenStudy (aaronandyson):

@Science_ALLY

OpenStudy (science_ally):

dude I cant help with this im only in algebra1 sowwy

OpenStudy (triciaal):

review sin 30 = cos 60 sin^2 + cos^2 = 1 tan = sin/cos

OpenStudy (aaronandyson):

now?

OpenStudy (triciaal):

2 sin^2(2A) = 1 - Cos 2A

OpenStudy (triciaal):

60 = 2* 30 sorry this stuff is not fresh in my mind does any of this give you any ideas?

OpenStudy (anonymous):

just put in values for sin 30 =1/2,sin 60 sq.rt3/2,tan 30 sq.rt 3/2 hence x =3.1

OpenStudy (aaronandyson):

Show your work.